OpenStudy (anonymous):

What is the point-slope equation of the line with slope that goes through the point (-2, 7)?

pooja195 (pooja195):

Whats the slope?

pooja195 (pooja195):

Is it given? Because without it you can't do this...

OpenStudy (anonymous):

5/2

pooja195 (pooja195):

ok so we use the point slope formula >:) \[\huge~\rm~y-y_1=m(x-x_1) \]

OpenStudy (anonymous):

yes

pooja195 (pooja195):

m=slope so in place of m we have 5/2 x1=-2 y1=7 can you substituite these values into the equation?

pooja195 (pooja195):

Need help?

pooja195 (pooja195):

you switched the x and y values try again

pooja195 (pooja195):

y-7=5/2(x+2) 2 negs make 1 pos